# A tale of two localizations: coexistence of flat bands and Anderson localization in a photonics-inspired amorphous system

### v1

@@ IV.2 Anderson-like localization in Class A
We stress that we cannot strictly rule out the possibility that an extended range of delocalized states for amorphous class A systems is a finite size effect. However, this observation certainly calls for future work, both numerically and analytically, to determine whether amorphous disorder can be fit into the AZ symmetry classification. We also note that the system sizes we simulate are on the scale of experimental topological photonics systems, thus even if delocalized states only appear so as a result of finite size, they may be effectively metallic for applications.
@@ V Conclusion
The methods we use to study localization in amorphous systems may also be applicable to studying newly-discovered fractional states in amorphous Chern insulators [42] and amorphous chiral spin liquids [43]. We also emphasize that our model system offers an approach to study introducing amorphousness perturbatively through combining local diagonal flips in the dual graph of a system.
@@
  • How do we understand the envelope structure observed in the effective localization spectrum in amorphous systems? The envelope does not appear to shrink with increasing system size, unlike in numerical simulations of onsite disorder. Furthermore, amorphous systems violate Mott’s argument: localized and delocalized states can exist at arbitrarily close energy spacing. We hypothesize that this is due to the discrete and constrained nature of amorphous disorder. Extending Mott’s argument to amorphous systems is an interesting avenue for further analytical and numerical exploration.
@@ A.1 Flat band structure
Numerical diagonalization of kagomé systems reveals that flat bands only exist when $\phi$ is a multiple of $\pi/6$. For the kagomé lattice, hexagons are the only secondary polygons and thus $6(2\phi+\pi)=2\pi m$ for $m\in\mathbb{Z}$ This condition indeed gives that $\phi$ must be a multiple of $\pi/6$ to observe a flat band. Note that flat band eigenstates never occur localized around the primitive triangles of the kagomé lattice because this geometry does not lead to destructive interference when the Hamiltonian acts on such a state.
@@ A.2 Flat band degeneracy
From Table 1, we see that the flat band degeneracy never exceeds $S/3+1$. This can be interpreted as the number of independent cycles that contain an even number of vertices (i.e. 0 or 2) of each primitive triangle. For the most part these cycles just come from secondary polygons, but we need to account for topology. We note that on a torus, the sum of cycles on all polygons vanishes, additionally, there are two non-contractible cycles. Overall, Eq. 7 gives:
$S/3-1+2=S/3+1.$ (8)
@@
The case of $\phi=0$: Recall, that the Hamiltonian Eq. 1 is given by a sum of terms supported on individual primitive triangles, each of which we will refer to as a local Hamiltonian. Each term has form $|1\rangle\langle 2|+|2\rangle\langle 3|+|3\rangle\langle 1|+h.c.$ with eigenvalues $2,-1,-1$. Each flat band state has energy $-2$ and this is only achievable if the state lies in the $-1$ subspace of the local Hamiltonian for each primitive triangle, since each site is acted upon by exactly two terms.
The $-1$ subspace is spanned by states that have amplitudes ±$\pm 1$ on the sites of any edge in the primitive triangle. Requiring that the amplitudes sum to zero on every primitive triangle leads us to conclude that the space of flat bands is equivalent to the space spanned by states supported on loops with the amplitude alternating ±$\pm 1$ along the loop. We note that a loop is permitted to go over the same edge multiple times, as occurs on the edges of Dirac strings in Type-II states introduced in Sec. III. Thus, a Type-II state is also considered a loop in this construction.
Finally, we can compute the number of such states on a kagomized graph with $S$ sites. If all secondary polygons are of even length, then the total flat band subspace spanned by the alternating states on each polygon plus the alternating states on non-contractible loops (their length can always be made even by making them use two edges of a triangle instead of one or visa versa). This is the case for periodic kagomé lattice and the number states coincides with Eq. (8): $S/3+1$. In the case when there are odd secondary polygons (they must come in pairs), we built a flat band state by combining any pair polygons pairwise if adjacent (Type-I states in Sec. III) or connecting them with a Dirac string otherwise (Type-II states in Sec. III). If there are $M$ odd secondary polygons, this procedure will yield $M-1$ linearly independent flat band states, leading to the degeneracy of $S/3$ confirmed by numerical simulations in Table 1.
@@
Figure 11: Top: Voronoi (dashed orange lines) diagram and kagomization (black) of a pinched graph that necessarily hosts Type-II flat band states. Bottom: Type-II flat band state on the kagomized graph. Blue (red) circles represent un-normalized state amplitude $+1$($-1$) and blue (red) hearts denote the Dirac string that connects the two disjoint secondary triangles, with amplitudes $+2$($-2$).
Appendix B Localization in amorphous model in zero-field
In Supplemental Material Fig. 3, we analyze amorphous systems with $\phi=0$, which belong to symmetry class AI. We find that states tend to localize near the upper tail of the single dispersive band. Near the fully delocalized state at $E=-4$, some states are also delocalized in character.
Figure 12: Scatter plot of the energy of every eigenstate of an $L_{0}=100$ system with $\phi=0$ and its effective localization length. At $E=-4$ is a completely delocalized state and at $E=2$ is the system flat band.

### v2

@@ IV.2 Anderson-like localization in Class A
In the following, we repeat the analysis of Sec. IV.1 for a system of Eq. (1) for a system with broken particle hole symmetry: in AZ class A. For subsequent calculations we set $\phi=4\pi/9$, such that the bulk remains gapped, see Fig. 7. The middle band has now broadened and the surrounding bands are no longer particle-hole symmetric. The upper band, however, retains qualitatively similar features to the class D case. In Fig. 12 in App. B we repeat the level spacings analysis of the previous section of subsets of this upper band and find qualitatively similar results to the analysis in class D. This agreement of localization properties between models in class A and class D with $E\neq 0$ was previously found and reasoned in a study of the thermal metal phase [50], however in those models, that did not involve amorphous disorder, all states were localized.
We stress that we cannot strictly rule out the possibility that an extended range of delocalized states for amorphous class A systems is a finite size effect. To evaluate this claim, one must numerically compute the conductance $g$ and mean free path $\ell$ for amorphous systems, compare $\ell$ to the system sizes we study, and $g$ to the criteria derived in [54]. We note, however, that the system sizes we simulate are on the scale of experiments with meta-materials, thus even if delocalized states only appear so as a result of finite size, they may be effectively metallic for applications.
IV.3 Localization-delocalization transition unique to amorphous systems
@@
The existence of this completely delocalized eigenstate is a consequence of the fixed coordination number in these graphs. For the uniform state, the local Hamiltonian appears the same on every site of the system and this state will occur in systems of all sizes. Remarkably, we observe that effective localization lengths of states nearby in energy to the fully delocalized state are significantly augmented, suggesting the existence of a localization-delocalization transition with no analog in generic disordered systems.
Other examples of delocalization at an isolated energy with a diverging localization length of localized states nearby are the 1D chain with random nearest-neighbor hopping and the 2D quantum Hall plateau transition, but in the present model the delocalized state arises as a result of fixed coordination number rather than symmetry (as in those cases). We generally expect that in models with a short-range Hamiltonian (such as Eq. 1) and fixed coordination number, a fully delocalized state and associated localization-delocalization transition exist if the energy of the delocalized state does not coincide with that of a flat band.
Figure 8: Top: Scatter plot of the energy of every eigenstate of an $L_{0}=64$ amorphous system with $\phi=0$ and its effective localization length. At E=4E=-4 is a completely delocalized state and at $E=2$ is the system flat band, colored in purple. Note that the localization length increases rapidly to the system size as $E\rightarrow-4$, and that the localization length is not uniquely defined for the massive degeneracy from flat band states at $E=2$. Bottom: Scatter plot of the energy of every eigenstate of an $L_{0}=64$ kagomé system with $\phi=0$ and onsite disorder of strength $W=2.5$, and its effective localization length.
As another example, in Fig. 8, we see this delocalization-localization in amorphous systems with $\phi=0$, which belong to symmetry class AI. Near the fully delocalized state at E=4E=-4, states become increasingly delocalized in character. We compare with systems with onsite-disorder but no amorphous disorder in Fig. 8. We set the disorder strength at $W=2.5$, adding a randomized energy to each site between [W/2,W/2][-W/2,W/2]. In the onsite disorder case, equivalent to the 2D Anderson model on a kagomé lattice, we see no fully delocalized state or localization transition.
The nature and critical properties of this localization-delocalization transition are interesting venues for future work. These observations, as well as the possibility of an extended range of metallic states in class A, also call for further study, both numerical and analytical, to expand the AZ symmetry classification to include amorphous disorder.
@@ V Conclusion
The methods we use to study localization in amorphous systems may also be applicable to studying newly-discovered fractional states in amorphous Chern insulators [56] and amorphous chiral spin liquids [57]. There is a considerable body of work on how flat bands in electronic systems modify interacting electron physics [58, 59, 60], which has expanded considerably with the advent of moire flat bands; while control of photon-photon interactions was one reason for experimental studies of amorphous photonic systems, further investigation of interactions in the photonic case is also needed. We also emphasize that our model system offers an approach to study introducing amorphousness perturbatively through combining local diagonal flips in the dual graph of a system.
@@
  • How do we understand the envelope structure observed in the effective localization spectrum in amorphous systems? How do we reconcile the coexistence of localized and delocalized states at arbitrarily close energies in Fig. 6 with Mott’s argument, which forbids such coexistence in generic disordered systems?
@@ A.1 Flat band structure
Numerical diagonalization of kagomé systems reveals that flat bands only exist when $\phi$ is a multiple of $\pi/6$. For the kagomé lattice, hexagons are the only secondary polygons and thus $6(2\phi+\pi)=2\pi m$ for $m\in\mathbb{Z}$ . This condition indeed gives that $\phi$ must be a multiple of $\pi/6$ to observe a flat band. Note that flat band eigenstates never occur localized around the primitive triangles of the kagomé lattice because this geometry does not lead to destructive interference when the Hamiltonian acts on such a state.
@@ A.2 Flat band degeneracy
From Table 1, we see that the flat band degeneracy never exceeds $S/3+1$. This can be interpreted as the number of independent cycles that contain an even number of vertices (i.e. 0 or 2) of each primitive triangle, we denote it as $C$. For the most part these cycles just come from secondary polygons, but we need to account for topology. We note that on a torus, the sum of cycles on all polygons vanishes, additionally, there are two non-contractible cycles. Overall, with Eq. 7, this yields
C=S/31+2=S/3+1.C=S/3-1+2=S/3+1. (8)
@@
The case of $\phi=0$: Recall, that the Hamiltonian Eq. 1 is given by a sum of terms supported on individual primitive triangles, each of which we will refer to as a local Hamiltonian. Each term has form $|1\rangle\langle 2|+|2\rangle\langle 3|+|3\rangle\langle 1|+h.c.$ with eigenvalues 2,1,12,-1,-1. Each flat band state has energy 2-2 and this is only achievable if the state lies in the 1-1 subspace of the local Hamiltonian for each primitive triangle, since each site is acted upon by exactly two terms.
The 1-1 subspace is spanned by states that have amplitudes $\pm 1$ on the sites of any edge in the primitive triangle. Requiring that the amplitudes sum to zero on every primitive triangle leads us to conclude that the space of flat bands is equivalent to the space spanned by states supported on loops with the amplitude alternating $\pm 1$ along the loop. We note that a loop is permitted to go over the same edge multiple times, as occurs on the edges of Dirac strings in Type-II states introduced in Sec. III. Thus, a Type-II state is also considered a loop in this construction.
Finally, we can compute the number of such states on a kagomized graph with $S$ sites. If all secondary polygons are of even length, then the total flat band subspace spanned by the alternating states on each polygon plus the alternating states on non-contractible loops (their length can always be made even by making them use two edges of a triangle instead of one or visa versa). This is the case for periodic kagomé lattice and the number states coincides with Eq. (8): $S/3+1$. In the case when there are odd secondary polygons (they must come in pairs), we built a flat band state by combining any pair polygons pairwise if adjacent (Type-I states in Sec. III) or connecting them with a Dirac string otherwise (Type-II states in Sec. III). If there are $M$ odd secondary polygons, this procedure will yield M1M-1 linearly independent flat band states, leading to the degeneracy of $S/3$ confirmed by numerical simulations in Table 1.
@@
Figure 10: Top: Voronoi (dashed orange lines) diagram and kagomization (black) of a pinched graph that necessarily hosts Type-II flat band states. Bottom: Type-II flat band state on the kagomized graph. Blue (red) circles represent un-normalized state amplitude +1+1(1-1) and blue (red) hearts denote the Dirac string that connects the two disjoint secondary triangles, with amplitudes +2+2(2-2).
Appendix B Nearest-level spacings distributions analyses
The set of nearest level spacings $\Delta E$} for a system is defined as the absolute difference between all nearest neighbors in an ordered set of the system’s eigenenergies. To define a distribution of the spacings, spacings are first normalized by the mean spacing to define $s=\Delta E/<\Delta E>$. In the limit of infinite system size, nearest-level spacings distributions in Anderson model disordered systems have been shown to fit either a Poisson law, $P_{P}(s)=1/\delta e^{-(s/\delta)}$, or the Wigner surmise for the orthogonal class ($\beta=1$), $P_{O}(s)=\frac{\pi}{2\delta}se^{-(\pi/4s^{2})}$ or a hybrid function, [64]. These three cases correspond to systems with only localized states, only delocalized (metallic) states, and systems at the metal-insulator transition.
An important difference from the Anderson model is that, for $\phi\neq 0$, systems break time reversal symmetry. Thus in our analysis we use the Wigner surmise for the unitary class (class A), $\beta=2$, $P_{W}(s)=\frac{32}{\pi^{2}}s^{2}e^{-4s^{2}/\pi}$ [65]. The Wigner surmise for class D coincides with that of class A for states at sufficiently high energy [66]. We define subsets of all eigenstates using cutoff $\tilde{\xi}/L<1/3$ to approximate which states are localized, shown for example in Fig. 6 in the main text. We examine the largest system size only. For each amorphous system of this size, we choose all energies within a range $R$. Then we select energies $E_{i,R}$} in this range and find the nearest level spacings $\Delta E_{i,R}$} and normalize this set by its mean. We then compute a histogram for the normalized spacings $\Delta E_{i,R}$}. Using the same histogram bins, we repeat this procedure for all realizations, then find the average value and error bar for each bin. For each range $R$, we fit a Poisson distribution and a Wigner distribution and then compute the $p$-value from the Kolmogorov-Smirnoff statistic $p$ for each fit. A large value $p\sim 1$ indicates that the data is likely to have come from this type of distribution whereas $p\sim 0$ indicates that the data comes from a different distribution. Results of these analyses are shown in Fig. 11, 12, 13
In amorphous systems with $\phi=0$, in symmetry class AI, we find that states tend to localize near the upper tail of the single dispersive band. In Supplemental Figure 13, we execute a nearest level-spacings analysis of the states suggested to localize, and confirm that these states are indeed localized. Note that since these systems are time-reversal-symmetric, we use the Wigner surmise for the orthogonal class ($\beta=1$), $P_{O}(s)=\frac{\pi}{2\delta}se^{-(\pi/4s^{2})}$. Our analysis suggests that states near the flat band at $E=2$ are localized. An intriguing question for future work is to determine whether adjacency to the flat band plays a role in localizing these states.
Figure 13: Nearest level-spacings distributions of states in the localized energy range for an amorphous systems of size $L_{0}=100$ with $\phi=0$. The localized energy range is set to $E_{loc}=(1.6,1.99)$, thus excluding the flat band at $E=2$. Best-fits for a Poisson law distribution and Wigner distribution are shown in orange and green, respectively. The distribution was drawn from the same distribution as the Poisson law.